Mechanics

A projectile is thrown with a velocity of 10underoot2 m/s at an angle of 45°w.r.t horizontal. The interval betn the moments when speed is underoot 125 m/a is?????

Anvita Mahajan
7 Years agoGrade

1 Answer

Rishab
7 Years ago
First take v=√125m/s then find hmax,
Hmax=u2sin2theta/2g
=(125*1/2)/20
=25/8
Now, we can find time
 
=T=√2h/g
Since,T=5/2s or 2.5s
 
=√2*25/8